The Kinetic Energy of a 0.5-kg monkey fist moving at 5m/s is


1
Expert's answer
2021-01-11T16:35:26-0500

By the definition of the kinetic energy, we have:


"KE=\dfrac{1}{2}mv^2=\dfrac{1}{2}\cdot 0.5\ kg\cdot(5.0\ \dfrac{m}{s})^2=6.25\ J."

Answer:

"KE=6.25\ J."